3Hg(OH)2 + 2H3PO4 → Hg3(PO4)2 + 6H2O


How many grams of mercury (II) hydroxide are required to produce 30.50 grams of water?



1
Expert's answer
2021-01-13T13:51:03-0500

moles of water 30.5g/18.0g = 1.69 mole

Moles of mercury (II) hydroxide will be equal to half moles of water

So moles of mercury (II) hydroxide = 1.69/2 = 0.845 moles

grams of mercury (II) hydroxide are required

= 0.845mole × 234.61g/mole

= 198.24g